OpenStudy (fools101):

Please help! Solve the following equation for x. 12x - 8 = 9x + 16 A) x=2 B) x=8 C) x=8/3 D) x=-8

OpenStudy (anonymous):

3x=24 x=8 the answer is B

OpenStudy (anonymous):

answer is B) x =8 since 12x-9x = 16+8 so 3x=24 and hence x =24/3=8

OpenStudy (fools101):

Thank u REMINDER and thank u JonnyMcA ^ ^

OpenStudy (anonymous):

B

OpenStudy (fools101):

Thank u Wangkan07 ^ ^

OpenStudy (anonymous):

youre welcome

OpenStudy (fools101):

^ ^...

OpenStudy (anonymous):

nice

OpenStudy (fools101):

Thanks! lol.